OpenStudy (anonymous):

use the formular nCr to evaluate the expression.8C5/4C3

OpenStudy (anonymous):

\[nCr = nPr/r! => n!/((n-r)!r!)\]

OpenStudy (anonymous):

I always go to pascals triangle

OpenStudy (anonymous):

8C5 = 8!/(3!5!) = (8*7*3!)/3! = 56 4C3 = 4!/(3!) = 4*3!/3! = 4 8C3/4C3 = 56/4 = 14

OpenStudy (anonymous):

And then how do you figure out the whole answer because isn't it suppose to come to one answer all toghether?

OpenStudy (anonymous):

because it's only suppose to be one answer because it's a fraction the 8c5 is on top and the 4c3 is on bottom

OpenStudy (anonymous):

1 4 6 4 1 <-- 4C3 = 4 1 5 10 10 5 1 1 6 15 20 15 6 1 1 7 21 35 35 21 7 1 1 8 28 56 70 56 28 8 1 <-- 8C5 = 56 56/4=14 Of course I would only list P's Triangle when I can do it faster than cyter's way

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

ur welcome!